OpenStudy (anonymous):

WILL FAN! Which expression has a value of 6, using only four 4s and any of the following: + , – , · , ÷, exponents, and/or parentheses? A. (4 + 4 + 4) ÷ 4 B. (4 · 4) – (4 + 4) C. (4 + 4) ÷ 4 + 4 D. 44 ÷ (4 · 4)

OpenStudy (mathstudent55):

You need to do all the choices with the correct order of operations.

OpenStudy (anonymous):

The answer is C. Using order of operations, it equates to (8) / 4 + 4. Because multiplication occurs before addition in PEMDAS order of operations, this becomes (8 / 4) + 4 = 2 + 4 = 6.

OpenStudy (mathstudent55):

In each one, first do any operations inside parentheses.

OpenStudy (anonymous):

it's not a because a is 3

OpenStudy (anonymous):

b is 16

OpenStudy (anonymous):

c is 6

OpenStudy (anonymous):

so its c